OpenStudy (anonymous):

Solve: 1+square root 3-x=x

OpenStudy (anonymous):

\[1+\sqrt{3-x}=x\]

OpenStudy (anonymous):

equal to nothing??

OpenStudy (anonymous):

the is the problem i was given

OpenStudy (anonymous):

\[1+\sqrt{3-x}=x\] \[\sqrt{3-x}=x-1\] by squaring both sides 3-x=x^2+1-2x 0=x^2+1-2x-3+x 0=x^2-x-2 or x^2-x-2=0 x^2-2x+x-2=0 x(x-2)+2(x-2)=0 (x-2)(x+2)=0 x=2 or x=-2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome can you give medal please

OpenStudy (anonymous):

thanks........................!!